OpenStudy (anonymous):

Evaluate arccot(10/23) in radians and round to 2 decimals. I got this wrong when I entered 1/tan-1(10/23)=2.44 what did i do wrong?

OpenStudy (anonymous):

\[\cot^{-1}(x) \neq \frac{1}{\tan^{-1}(x)} \]

OpenStudy (anonymous):

that's what I used i think but the professor said my answer was wrong?

OpenStudy (anonymous):

To see this, let \[y = \cot^{-1}(x) \rightarrow \cot(y) = x\] Then \[\tan(y) = \frac{1}{x} \] and so \[ y = \tan^{-1}\left(\frac{1}{x}\right) \] Which means that \[ \cot^{-1}(x) = \tan^{-1}\left(\frac{1}{x}\right) \]

OpenStudy (anonymous):

No, it's NOT equal. The slash through the equals sign means not equal.

OpenStudy (anonymous):

I'm sorry, I'm tired. so my question would be tan^-1(1/(10/23)) ???

OpenStudy (anonymous):

Yes. While it's true that \[\cot(x) = \frac{1}{\tan(x)} \] the same is NOT true for the inverse functions. Instead, as shown above, \[\cot^{-1}(x) = \tan^{-1}\left(\frac{1}{x}\right) \] so \[\cot^{-1}\left(\frac{10}{23}\right) = \tan^{-1}\left(\frac{23}{10}\right) \approx 1.16 \]

OpenStudy (anonymous):

oh, thank you so much! I kept getting these problems wrong as one person had me working them the other way.

OpenStudy (anonymous):

I am sorry to hear that. But, I suppose everybody makes mistakes, even people who are trying to help :)

OpenStudy (anonymous):

it makes sense now. thank you!